# AI-coupled HPC Workflow Applications, Middleware and Performance

### v1

@@ 2 Execution Motifs
There have been a variety of HyperParameter Optimization (HPO) approaches in the recent years for designing hyperparameters for deep learning on HPC, including using Bayesian optimization [59], evolutionary algorithms [60], or genetic algorithms [61]. These methods have been used in HPC for the development of AI surrogate models and traditional semi-analytic reduced models (e.g., to build an AI model predictive control for optimal real-time control landing of spacecraft [62]) or for building AI-based analysis routines (e.g., the MENNDL HPO is used to build models for cancer pathology to classify tumor-infiltrating lymphocytes [60]). Other examples of this motif include Neural Architecture Search (NAS) [63, 64], Large Language Model (LLM) searches [65, 66], Adaptive LLM training [67, 68]. Enabling dynamic interactions between learning agents like HPO, NAS or LLMs, and HPC simulations (with or without human intervention) is becoming increasingly used in several fields: computational fluids dynamics [69], fusion [20], energy physics [21], and climate science [70]. These studies combine HPC with AI to optimize the training process by steering the simulation to generate real-time data. The FASTRAN-IPS [20] workflow is an example of the training being done online in a strongly coupled fashion. The simulation generates data iteratively that is then used to train a surrogate model. At every step, the AI algorithm identifies gaps in the parameter space and steers the simulation to generate the missing data. Coupling AI and simulation with visualization is also used to allow users to interactively explore complex parameter spaces given by a large number of simulation runs and interactively steer or span new simulations to cover the missing parts in the parameter space. This is when reduced-form simulations developed through machine learning techniques are coupled with visualization to allow interactive feedback and steer the simulation to generate new data [21]. Estimates of the results of the full simulations at a fraction of the computational cost are used to steer the learning process, as the workflow requires a human to inspect the outcome of the simulation.
@@
The motifs presented in this section use different coupling and interaction patterns when executing HPC and AI for different scopes. These motifs are not mutually exclusive; we expect future large-scale workflows to combine several motifs. One such example deploys an AI-driven workflow, combining three motifs [71]: (1) it adaptively steer MD simulation ensembles; (2) it further couples this workflow with AI-based methods in a multistage pipeline fashion; (3) it uses adaptive execution to learn which parts of the parameter space have been sampled sufficiently and to trigger simulations from under-sampled regions. Each component depends on a distinct solution: the steering uses DeepDriveMD, the learning employs a 3D PointNet-based adversarial autoencoder [72], and the overall integration relies on a custom solution.
Scientific computing infrastructure trends toward an integrated research infrastructure (IRI) [73] that will unite data resources, experimental user facilities, and advanced computing resources to accelerate the deployment of large-scale simulations and workflows. Such integrated infrastructures will accelerate the trend of coupling multiple HPC simulations, surrogate models, and digital twins with AI analysis at multiple levels.
@@ 3 AI-coupled HPC Workflow Management Frameworks
For two decades, automation, simplification, and workflow management optimization have been active research topics in the HPC and distributed systems communities. This has led to developing many libraries, frameworks, and tools for domain scientists to advance their scientific agenda. More recently, diverse research communities in both industry and academia have been working on allowing AI frameworks to go beyond being a standalone black-box operator to facilitate their inclusion into more complex workflows. While the previous section highlighted that several studies involving AI-coupled HPC workflows relied on custom solutions with low portability and high maintenance cost, we observe a convergence of the efforts related to workflow management in both the HPC and AI communities with the recent development of workflow frameworks specifically targeting AI-coupled HPC workflows. We review some of these frameworks and identify their connections with the execution motifs identified in the previous section. Then, we identify several challenges related to the development of those frameworks.
@@ 3.1 Overview of Workflow Management Frameworks
Table 2 lists these frameworks and summarizes their main characteristics. When applicable, we also indicate the underlying workflow technology on which the framework is built. We distinguish frameworks developed with a specific application domain from more general-purpose tools that address issues shared by several application domains, e.g., optimizing data access patterns and related I/O operations.
@@
CANDLE [74, 75] Cancer research Steering — Multistage Pipeline — Adaptive Training SWIFT/T
DeepDriveMD [76] Biophysics/Molecular Simulations Steering — Multistage Pipeline — Adaptive Training Radical
MuMMI [77] Biophysics/Molecular Simulations Steering — Multistage Pipeline — Adaptive Training Maestro [78]
@@
Colmena [23] General-purpose steering of ensembles Steering — Adaptive Training Parsl
SmartSim [80, 81] A scalable, open-source, multi-physics simulation API. Steering — Multistage Pipeline — Digital Replica — Adaptive Training Redis
@@
The DeepDriveMD [76] framework has been designed to effectively reduce the dimensionality of MD simulations to get insights into how proteins fold. The objectives are to build latent representations that automatically correspond to biophysically relevant reaction coordinates (RCs) and drive MD simulations to automatically sample potentially novel conformational states based on these RCs. The AI-coupled HPC workflow underlying DeepDriveMD is composed of four main steps: (i) A large ensemble of MD simulations generates the initial data needed to (ii) train a Deep Learning algorithm, (iii) whose inference step identifies new starting points in the ensemble of MD simulations, and (iv) enable the actual steering of the ensemble. This AI-driven steering may consist of either expanding the pool of initial MD simulations or pruning simulations that seem to be stuck in meta-stable states.
The objective of the Multiscale Machine-Learned Modeling Infrastructure (MuMMI) [77] is to couple a large-scale parallel simulation of a macro-scale model based on dynamic density functional theory that spans micrometer length- and millisecond time-scales with a micro-scale model based on high-fidelity MD simulations. MuMMI completes this workflow with an AI-driven importance-sampling framework to bridge the resolution gap between the models. It identifies sub-regions of scientific importance in the results of the macro-scale model, which then serves as candidate inputs to many thousands of coarse-grained MD simulations. The results of these fast-responding simulations are analyzed in situ to avoid the cost of storing them on disk and feed a self-healing feedback loop to improve the parameters of the macro model. To manage such an AI-coupled HPC workflow and steer the framework towards the highest scientific interest simulations, MuMMI relies on the Maestro workflow conductor [78]. Regarding data management, the coarse-grained MD simulations launched by MuMMI could generate PB of data in over 400 million files, a number to multiply by 10 when AI is added to the workflow. MuMMI implements several local in situ and data aggregation strategies to reduce the demand on the filesystem.
The IMPECCABLE [79] project aims at defining an integrated modeling pipeline to identify effective chemical compounds to fight COVID-19, by screening tens of billions of molecules. IMPECCABLE uses AI methods to improve both the initial sampling of candidate molecules and the design of surrogate models of multi-physics simulations. It also integrates these AI components and HPC simulations into a single unified three-stage pipeline. IMPECCABLE relies on the Radical-Cybertools [84] to manage the execution of this heterogeneous AI-coupled HPC workflow on leadership computing facilities.
In addition to these frameworks proposed to address the challenges related to the coupling of AI and HPC in a specific scientific domain, we detail hereafter, the intents and characteristics of more general-purpose frameworks that propose solutions applicable to several domains.
Colmena [23] is a framework allowing users to steer simulation ensembles using an AI model. It handles the dispatch of highly computational simulation runs and model retraining tasks, collating the simulation results and invocating the AI model. It relies on the Parsl workflow management system to handle the heavy computations and their results. Colmena also offers a framework called the Thinker to design a decision-making multi-agent process adapted to the users’ needs. Each agent can post task requests, handle results, or react to resource-related events using Redis queues. The AI component thus serves as the actual pilot of the simulation campaigns, deciding which inputs to select, in what order, and when to execute each simulation run.
The SmartSim [80, 81] framework, developed by Hewlett Packard Enterprise enables the composition of HPC simulations and AI models into an in situ workflow. The HPC simulations act as data producers, that generate training and inference data while the AI component consumes this data either for training or inference. SmartSim relies on an in-memory database to store any information shared by the HPC simulations and the AI components for the duration of the run, which allows for interactions without having to pay for expensive file-based accesses on a shared parallel file system. It also exposes a multi-language communication API based on the publish/subscribe paradigm, which facilitates the connection of scientific codes, usually written in C, C++, or Fortran, and the Python-based ML frameworks.
Stimulus [82] is a data management library designed to ingest scientific data into popular AI frameworks efficiently and safely. It aims to address the problem of data format and access pattern incompatibility when HPC simulations and AI components are combined in one of the motifs introduced in the previous section. Indeed, following a strongly coupled approach in which HPC simulations and AI components would share the same format is intractable and will duplicate efforts. The approach followed by Stimulus is to unify several popular scientific data formats (e.g., NetCDF, HDF5, ADIOS) under a single interface that hides the complexity of each format. Additionally, Stimulus proposes a generic data ingestion pipeline that any tensor-operator-based AI framework can execute.
NoPFS (Near-Optimal PreFetching System) [83] is Machine Learning I/O middleware aims to reduce shared filesystem contention when ingesting data to train a model. To this end, NoPFS leverages the clairvoyance an I/O system can have by determining how access will be made from the initial seed of a pseudo-random number generator. This clairvoyance allows NoPFS to efficiently prefetch data, exploiting the whole storage hierarchy available on a compute node and minimizing access to the shared filesystem. It is not an AI-coupled HPC workflow system per se, but using NoPFS to manage the I/O model training would free some throughput on the parallel file system, which could then be used by HPC simulations.
Note that the integration of AI components into more complex analysis workflows is not limited to the HPC community. For instance, the Kafka-ML project [85] aims to connect the Apache Kakfa distributed messaging system to AI models to get more insight from the continuous data streams from the Internet of Things sensors. Similarly, the Redis NoSQL database server includes a module, RedisAI, allowing to develop inference servers fed with Redis data streams [81] In the context of initiatives such as the Integrated Research Infrastructure from DOE that aims to connect experimental facilities producing such data streams to leadership-class HPC systems, we envision that stream processing frameworks will become part of complex AI-coupled HPC workflows soon.
@@ 3.2 Challenges
In this section, we considered frameworks developed to automate the execution of AI-coupled HPC workflows in different scientific domains or designed to address some common challenges from the perspective of their matching with the proposed execution motifs. However, the sheer amount of data handled by these frameworks (e.g., ”scaling to full memory ($>$ PB) on the largest supercomputers” [74]), the high level of task concurrency they manage (e.g., 16k concurrent tasks for DeepDriveMD [76]) or both (e.g., 36k concurrent tasks and several PB of raw data for MuMMI [77]) can have some important performance implications. In the next section, we focus on these performance considerations of AI-coupled HPC workflows.
@@ 4 Performance
As we delve deeper into the intricacies of AI-coupled HPC workflows, it becomes evident that several open issues emerge, particularly regarding interoperability, data handling, and resource allocation. These challenges underscore the importance of understanding the performance and highlight broader opportunities to enhance the efficiency and effectiveness of such workflows.
@@ 4.1 Workflow Components
Each workflow component has performance implications individually and in their coupling. In general, for AI, especially for AI raining, there is a trade-off between computation vs. communication [87]. Offloading to AI accelerators requires data movement and may be largely affected by how the accelerators are integrated into the main supercomputer. GPUs typically achieve lower latencies for large batch queries, although they incur significant loading overhead; conversely, CPUs usually exhibit lower loading latencies and excel with smaller batch sizes [88]. This topic has been studied comprehensively in Ref. [89] for an AI-driven simulation across cross-facility infrastructure performing simulations on a traditional HPC supercomputer while offloading the AI model to a Cerebras accelerator, a wafer-scale application-specific integrated circuit (ASIC) device.
@@
(a) Machine learning training and deployment pipeline
@@
(b) Pipeline for typical HPC applications
@@
The HPC pipeline illustrated in Figure 6(b) has typical I/O and communication network performance considerations. As HPC simulations evolved to include GPU kernels and/or AI in their codes, the trade-offs described in the previous paragraph are also becoming relevant for the HPC components. When deployed together, the AI and HPC components will overlap different stages of their respective pipelines and face new performance bottlenecks that are unnecessary during individual runs. For example, the Data Production phase of the HPC application will typically overlap with Data Ingestion for the Steering motif, which might cause concurrent read/write access to the same datasets, resulting in performance penalties due to I/O contention.
@@ 4.2 Coupling Approaches and Protocols
The critical insight to remember is that the performance of a specific workflow is substantially influenced not only by the efficiency of its elements but also by the way those elements are interconnected and integrated. Several studies have attempted to investigate how the coupling of workflow components affects overall performance [100, 101, 102, 103, 104, 105, 106]. Several studies have attempted to investigate how the coupling of workflow components affects overall performance [100, 101, 102, 103, 104, 105, 106].
@@ 4.3 Motif Performance
Steering In-situ coupling, loose or tight coupling strategies Data transfer between simulation and AI, latency in decision-making
@@
Table 3: Summary of execution motif characteristics, performance bottlenecks, and coupling strategies [107].
4.3.1 AI-based steering ensembles of simulations
The Steering motif uses machine-learned models to steer ensembles of simulations. Colmena is an example of such a framework for implementing such a task on HPC and has been demonstrated for molecular design [23]. This type of architecture is similar to distributed reinforcement learning applications, designed to maximize a reward function by running many simulations (environments) in parallel [108]. Dataflow performance is the main challenge for these workflows, especially when a centralized server for task planning that can become a performance bottleneck.
In the Steering motif, data is ingested from parallel simulations, processed, and used to train machine learning models. These models are then evaluated for effectiveness in guiding simulations and deployed in an HPC environment to steer ensembles. For example, Colmena has three processes: a thinker, a task server, and a worker. The thinker communicates with the task server using a request-response architecture, specifically Redis queues, which use the Redis Serialization Protocol (RESP). Workers are scaled up on the system using Parsl [109].

### v2

@@ 2 Execution Motifs
There have been numerous studies applying Hyperparameter Optimization (HPO) techniques–including Bayesian optimization [58], evolutionary strategies [59], and genetic algorithms [60]–to improve surrogate models and reduced-order representations for HPC applications. These approaches have been used across domains such as cancer pathology [59], spacecraft landing control [61], and molecular simulations. Similarly, Neural Architecture Search (NAS) [62, 63] and Large Language Model (LLM) training [64, 65, 66] increasingly leverage simulation-aware feedback loops to explore model configurations more efficiently. Dynamic coupling of these learning agents with HPC simulations–whether automated or human-guided–is gaining traction in fields such as fusion [19], fluid dynamics [67], energy systems [20], and climate science [68]. The FASTRAN-IPS workflow [19] exemplifies this approach by training a surrogate model iteratively while steering simulations to explore undersampled regions of the parameter space. AI is also paired with visualization systems to allow users to interactively explore results and launch new simulations as needed [20].
@@
The motifs presented in this section use different coupling and interaction patterns when executing HPC and AI for different scopes. These motifs are not mutually exclusive; we expect future large-scale workflows to combine several motifs. One such example deploys an AI-coupled HPC workflow, combining three motifs [69]: (1) it adaptively steer MD simulation ensembles; (2) it further couples this workflow with AI-based methods in a multistage pipeline fashion; (3) it uses adaptive execution to learn which parts of the parameter space have been sampled sufficiently and to trigger simulations from under-sampled regions. Each component depends on a distinct solution: the steering uses DeepDriveMD, the learning employs a 3D PointNet-based adversarial autoencoder [70], and the overall integration relies on a custom solution.
Scientific computing is evolving towards integrated research infrastructures (like IRI [71]) that will unite data resources, experimental user facilities, and advanced computing resources to accelerate the deployment of large-scale simulations and workflows. Such integrated infrastructures will accelerate the trend of coupling multiple HPC simulations, surrogate models, and digital twins with AI analysis at multiple levels.
@@ 3 AI-coupled HPC Workflow Management Frameworks
For two decades, automation, simplification, and workflow management optimization have been active research topics in the HPC and distributed systems communities. This has led to developing many libraries, frameworks, and tools for domain scientists to advance their scientific agenda. More recently, diverse research communities in both industry and academia have been working on allowing AI frameworks to go beyond being a standalone black-box operator to facilitate their inclusion into more complex workflows. While the previous section highlighted that several studies involving AI-coupled HPC workflows relied on custom solutions with low portability and high maintenance cost, we observe a convergence of the efforts related to workflow management in both the HPC and AI communities with the recent development of workflow frameworks specifically targeting such workflows. We review some of these frameworks and identify their connections with the execution motifs identified in the previous section. Then, we identify several challenges related to the development of those frameworks.
@@ 3.1 Overview of Workflow Management Frameworks
To complement the execution motifs introduced earlier, Table 2 summarizes key frameworks developed for AI-coupled HPC workflows. Some are domain-specific, while others offer general-purpose capabilities. In the discussion below, we cluster frameworks with similar architectural patterns and highlight representative use cases from each category.
@@
CANDLE [72, 73] Cancer research Dynamic Orchestration — Multistage Pipeline — Adaptive Training SWIFT/T
DeepDriveMD [74] Biophysics/Molecular Simulations Dynamic Orchestration — Multistage Pipeline — Adaptive Training Radical
MuMMI [75] Biophysics/Molecular Simulations Dynamic Orchestration — Multistage Pipeline — Adaptive Training Maestro [76]
@@
Colmena [22] General-purpose steering of ensembles Dynamic Orchestration — Adaptive Training Parsl
EXARL[78, 79] General-purposeReinforcement Learning control Dynamic Orchestration — Adaptive Training MPI
GlobusCompute [80, 81] Function-as-a-Service on federated resources framework Multistage Pipeline — Distributed Models Globus, Parsl
INTERSECT [82] Interconnected Science Ecosystem Distributed Models — Dynamic Orchestration — Multistage Pipeline — Digital Replica -
SmartSim [83, 84] A scalable, open-source, multi-physics simulation API. Dynamic Orchestration — Multistage Pipeline — Digital Replica — Adaptive Training Redis
@@
Several domain-specific frameworks–DeepDriveMD, MuMMI, and IMPECCABLE–share a common loop of simulation, AI-based analysis, and adaptive steering. These frameworks often operate on ensembles of simulations, use AI to select promising subregions or filter states, and incorporate feedback loops to optimize sampling or model fidelity. DeepDriveMD [74] focuses on protein folding. It uses deep learning to construct latent representations of molecular conformations and steers ensembles of molecular dynamics (MD) simulations accordingly. RADICAL-Cybertools orchestrate simulation workflows and model-driven steering. MuMMI [75] couples macro-scale density functional theory simulations with micro-scale MD simulations. An AI-driven importance-sampling engine bridges the two scales, dynamically identifying sub-regions for follow-up simulation. MuMMI employs the Maestro workflow system and in situ data aggregation to manage the high data volume. IMPECCABLE [77] integrates surrogate modeling and AI-guided sampling into a three-stage pipeline for drug discovery. It uses Radical-Cybertools to manage the workflow across leadership-class systems.
Beyond domain-specific frameworks, a variety of general-purpose workflow systems have been developed to support AI-HPC integration across scientific fields. These frameworks differ in orchestration models, data handling strategies, and support for federated or in situ execution.
Colmena [22] supports AI-guided steering of simulation ensembles. Built on the Parsl workflow system, Colmena enables concurrent execution of simulations, AI retraining, and decision-making agents. It introduces the “Thinker” interface, which coordinates task scheduling using Redis-backed queues and allows for flexible, user-defined control strategies.
EXARL [78, 79] enables reinforcement learning at HPC scale by implementing a learner/actor architecture within a parallel MPI-based training loop. It supports concurrent learning and simulation with OpenAI Gym environments and has been integrated with domain-specific workflows such as CANDLE for cancer research.
Globus Compute [80, 81] (formerly FuncX) is a federated Function-as-a-Service (FaaS) platform designed for cross-site workflows. It uses a REST-based model to invoke user-defined functions across distributed computing endpoints, including edge devices, clusters, and HPC systems. It has been demonstrated in real-time AI analysis pipelines for experimental data, such as x-ray ptychography.
INTERSECT [82] is a distributed, microservices-based framework developed at Oak Ridge National Laboratory to connect robotic labs, instruments, and compute facilities. It enables self-driving labs through AI-guided experiments and real-time data streaming. The system includes a control message queue and scientific data layer, and has been applied in electrochemistry experiments.
SmartSim [83, 84] is an in situ execution framework that couples simulations with AI models using an in-memory Redis database. It supports data exchange between Fortran/C++ simulation codes and Python-based ML tools via a publish/subscribe model. This design allows low-latency communication without reliance on file-based I/O.
Stimulus [85] addresses data ingestion challenges by providing a unified interface across scientific data formats (e.g., HDF5, NetCDF, ADIOS) for use in AI frameworks such as PyTorch and TensorFlow. It simplifies access pattern mismatches and decouples HPC simulation output from AI pre-processing code.
NoPFS [86] is a Machine Learning I/O middleware that optimizes data prefetching during model training. By leveraging knowledge of randomized data access patterns, it reduces contention on shared filesystems. Though not a full workflow system, it improves training throughput and indirectly benefits simulation workloads running on the same infrastructure.
AI integration into real-time and streaming workflows extends beyond HPC contexts. For instance, Kafka-ML [87] connects Apache Kafka with AI models to process data from IoT sensors. Similarly, RedisAI enables inference directly within Redis streams [84]. As efforts like the DOE Integrated Research Infrastructure expand, we expect stream processing frameworks to play a central role in AI-coupled HPC workflows.
@@ 3.2 Challenges
In this section, we considered frameworks developed to automate the execution of AI-integrated workflows in different scientific domains or designed to address some common challenges from the perspective of their matching with the proposed execution motifs. However, the sheer amount of data handled by these frameworks (e.g., ”scaling to full memory ($>$ PB) on the largest supercomputers” [72]), the high level of task concurrency they manage (e.g., 16k concurrent tasks for DeepDriveMD [74]) or both (e.g., 36k concurrent tasks and several PB of raw data for MuMMI [75]) can have some important performance implications. In the next section, we focus on these performance considerations of such workflows.
@@ 4 Performance
As we delve deeper into the intricacies of AI-integrated workflows, it becomes evident that several open issues emerge, particularly regarding interoperability, data handling, and resource allocation. These challenges underscore the importance of understanding the performance and highlight broader opportunities to enhance the efficiency and effectiveness of such workflows.
@@ 4.1 Workflow Components
Each workflow component has performance implications individually and in their coupling. In general, for AI, especially for AI training, there is a trade-off between computation vs. communication [89]. Offloading to AI accelerators requires data movement and may be largely affected by how the accelerators are integrated into the main supercomputer. GPUs typically achieve lower latencies for large batch queries, although they incur significant loading overhead; conversely, CPUs usually exhibit lower loading latencies and excel with smaller batch sizes [90]. This topic has been studied comprehensively in Ref. [91] for an AI-driven simulation across cross-facility infrastructure performing simulations on a traditional HPC supercomputer while offloading the AI model to a Cerebras accelerator, a wafer-scale application-specific integrated circuit (ASIC) device.
@@
((a)) Machine learning training and deployment pipeline
@@
((b)) Pipeline for typical HPC applications
@@
The HPC pipeline illustrated in Figure 7(b) has typical I/O and communication network performance considerations. As HPC simulations evolved to include GPU kernels and/or AI in their codes, the trade-offs described in the previous paragraph are also becoming relevant for the HPC components. When deployed together, the AI and HPC components will overlap different stages of their respective pipelines and face new performance bottlenecks that are unnecessary during individual runs. For example, the Data Production phase of the HPC application will typically overlap with Data Ingestion for the Dynamic Orchestration motif, which might cause concurrent read/write access to the same datasets, resulting in performance penalties due to I/O contention.
@@ 4.1.4 Model Deployment
4.1.5 Simulation
Traditional scientific HPC applications, such as computational fluid dynamics (CFD), or in general finite different, or finite element simulations, require handling the challenges of sustaining performance at scale both at the intra-node (OpenMP) and inter-node (MPI) level, as well as optimizing for on/off loading to the GPU (CUDA/HIP) [102, 103, 104]. Efficient use of hardware also requires handling the challenges of compute-bound versus memory-bound applications. There are performance challenges stemming from inter-process communication involving both collective operations and halo exchanges. On the GPU side, depending on the architecture there may be performance implications whether the network adapters are bound to the GPU directly for direct GPU-GPU communication across nodes (e.g., RDMA) and also based on whether the GPU has to access node RAM or has its own memory.
4.1.6 Data production
Data production involves the generation and storage of large datasets produced by the simulations. On current, large-scale HPC platforms, data moving from compute nodes to storage pass through several layers in a distributed hierarchy of resources (e.g., memory, burst buffers, I/O nodes, filesystem servers). The first critical bottleneck is due to file I/O of such datasets, which may be mitigated, depending on the access pattern, by efficient parallel I/O implementations. At the lower level, most systems have a parallel file system such as Lustre or GPFS, on top of which additional I/O libraries may be exploited, such as ADIOS2, HDF5, and netCDF. Current I/O libraries encounter challenges when staging data between HPC and AI due to their contrasting data access needs. HPC prioritizes high-bandwidth, often write-optimized, I/O for substantial large transfers across parallel file systems. Conversely, AI/ML workloads demand fast, random access to numerous small data samples for training, emphasizing read efficiency. This fundamental conflict in optimization-high-bandwidth sequential (HPC) versus low-latency random (AI)-makes current libraries less efficient in serving both application types optimally during data staging.
@@ 4.2 Coupling Approaches and Protocols
The critical insight to remember is that the performance of a specific workflow is substantially influenced not only by the efficiency of its elements but also by the way those elements are interconnected and integrated. Several studies have attempted to investigate how the coupling of workflow components affects overall performance [105, 106, 107, 108, 109, 110, 111].
@@ 4.3 Motif Performance
Dynamic Orchestration In-situ coupling, loose or tight coupling strategies Data transfer between simulation and AI, latency in decision-making
@@
Table 3: Summary of execution motif characteristics, performance bottlenecks, and coupling strategies.
4.3.1 Dynamic Orchestration
The Dynamic Orchestration motif uses machine-learned models to steer ensembles of simulations. Colmena is an example of such a framework for implementing such a task on HPC and has been demonstrated for molecular design [22]. This type of architecture is similar to distributed reinforcement learning applications, designed to maximize a reward function by running many simulations (environments) in parallel [112]. Dataflow performance is the main challenge for these workflows, especially with a centralized server for task planning that can become a performance bottleneck.
In the Dynamic Orchestration motif, data is ingested from parallel simulations, processed, and used to train machine learning models. These models are then evaluated for effectiveness in guiding simulations and deployed in an HPC environment to steer ensembles. For example, Colmena has three processes: a thinker, a task server, and a worker. The thinker communicates with the task server using a request-response architecture, specifically Redis queues, which use the Redis Serialization Protocol (RESP). Workers are scaled up on the system using Parsl [113].